# CosForce: A Force-Based General Pedestrian Model with Anticipation and Reaction Mechanisms

### v3

@@ 5.3 Discussion
In sections 5.1 and 5.2, based on simulations, we discussed the "catfish effect" and the process of crowd collapse, and provided an explanation based on the processes of critical phase transition. From the simulation results, we identified several densities of tipping point for crowd phase transitions, specifically $\rho=$ 9.8 and 10.4 ped/m2. These values are applicable only within the context of this simulation. Human body size varies significantly by gender, age, and other physiological factors, density is an insufficient metric. More importantly, it is essential to understand the critical conditions under which individuals lose control of their motion due to contact forces. All insights in section 5 were derived retrospectively from the perspective of simulation with inherently limitations and potential errors. Based on the presented results, several issues need to be discussed in this section.
The simplification of crowds into granular media interactions or even continuous fluids has essentially become a standard procedure in crowd analysis. Based on such simplifications, we need discuss how extent which crowds exhibit similarities to granular flows. In terms of similarities, crowds can demonstrate many characteristics of granular flows, such as the phenomenon of cross-channel formation, which exhibits solid-liquid flow-like interactions, as well as rheology and vibration compaction effects, typically observed in granular media. Considering the differences, the most fundamental distinction lies in the fact that crowds are self-propelled, and the motion processes involve internal energy, which is not the case in granular flows. Additionally, crowds are largely heterogeneous, with this heterogeneity not only evident in size and speed but also in the individual differences in psychological decision.
@@
Potential intentional pushing was reported during the Itaewon crowd crush incident (Kim & Kim, 2022). Similar aggressive behaviors, as manifestations of crowd heterogeneity, require attention in simulations. This is one of the purposes of establishing section 5.1, where we introduced dynamic pedestrians in the simulation to analyze relevant differences. In this consideration, another question raised: whether dynamic (or aggressive) pedestrians would accelerate the process of crowd collapse. In our simulation, introduction of dynamic pedestrians significantly destabilizes the state of the crowd. From the perspectives of mobility (average normalized speed) and symmetry (order parameter), each dynamic pedestrian acts like a microscopic magnet that induces local polarization in the crowd, as demonstrated in section 5.1. This manifests as an increase in observed dynamism and a decrease in symmetry, leading to sharp fluctuations in speed and density within local regions. Fig.17 and Fig.19 report that the equilibrium flow configuration (bi and four direction) in the simulation with dynamic pedestrians exhibits a lower order parameter. However, based on the velocity fields in Fig.16 and Fig.18, it can be concluded that this results from the statistical cancellation of polarizing effects.
@@ 6 Advantages and limitations
We have conducted extensive empirical validation and testing of the model. In this part, The most basic properties of the model will be discussed: how efficient the model is and under what circumstances the model may become ineffective.
@@ 6.1 Property of linear time complexity
      for $\textit{round}=1$ to $2$ do
[…]
@@
[…]
@@
                  case $(0,-1)$ do
                         direction_search = $(-1,0)$
@@
Benchmark tests were conducted on a laptop (Intel Core i7-11800H CPU, 16 GB memory) to evaluate simulation performance. In the simulation, we evaluated five configurations for the structured spatial basic units: an adaptive length and four fixed lengths (0.5 m, 1 m, 1.5 m, and 2 m). The setup involved a 100m × 100m (10,000 m²) simulation space in which agents were randomly generated with equal increments (2,000 pedestrians per run) over 3,000 iterations. The runtime results for each basic unit length configuration are presented in Fig.24.
@@
Fig.24 clearly demonstrate the linear time complexity characteristics of the model. The length of basic unit and runtime show statistically significant linear relationships. Comparisons across different groups reveal varying performance for different unit sizes. The overall trend indicates that larger unit sizes lead to lower computational efficiency. This trend is clearly supported by the linear fitting results. For unit lengths of adaptive size, 0.5m, 1m, 1.5m, and 2m, the fitted slopes are 0.37, 0.37, 0.40, 0.53, and 0.69, respectively.
@@ 6.2 Limitations
We also explored the limitations of the model, which mainly include two aspects: Intractability in task related with decision process. As shown in Fig.25 (a) and (b), even simulating some local tactical-level behaviors, such as short-term route choice, is beyond the capability of the model, which is also a common limitation of most operational-level models. Distortion in response to static environments. In certain static environments, pedestrians may become trapped even when space is available. Within the framework of force-based models, there exist equilibrium positions in static environments that cause the pedestrian’s force balance, resembling a black hole, where pedestrians passing nearby will drawn towards these equilibrium positions.
@@ 7 Conclusions
Empirical validation of the model was conducted for scenarios involving single-file motion and unidirectional flow. In simulations, we examined the phase separation process in crowds, including lane formation, stripe formation, and cross-channel formation, typically regarded as the phenomena of pedestrian self-organization. By analyzing entropy decrease and steady-state characteristics, quantitative observations clearly reveal the self-organization process. Furthermore, through mass gathering simulations, we quantitatively demonstrated the "catfish effect", a theoretical concept that currently lacks empirical validation. In this context, we analyzed polarization and the process of symmetry-breaking in crowd dynamics.
Finally, focusing on crowd collapse, we explored potential mechanisms through time series and flow-field analysis. Applying critical transition theory, we characterized crowd collapse as a phase transition process quantified by the order parameter. As crowd density increases, the first phase transition (second-order) occurs when the crowd becomes tightly compressed. At this stage, pedestrian interactions shift from self-driven constraints of fundamental diagrams to granular like dynamics. The second phase transition (first-order) occurs during crowd collapse. At this stage, further compression approaches the limit, elevating the order parameter. When the crowd reaches a critical state (characterized by high density and a persistently rising order parameter), collapse initiates.
Given the model’s simple rules and evaluation based on basic parameters (e.g., average normalized speed, order parameter), assessments of crowd dynamics can be intuitively derived, thus providing valuable insights and inspiration. However, since the findings are derived from the hindsight of simulation, their inherent limitations should not be overlooked.
@@ Appendix A Experiments
The empirical validation in section 3 was conducted using data from two sets of experiments: single-file motion and unidirectional flow. The datasets are available at: https://ped.fz-juelich.de/da/doku.php?id=start#data_section (Pedestrian Dynamics Data Archive). Detailed experimental configurations and procedural specifics can be found in the works of Cao et al. (2016) and Cao et al. (2017), as referenced.
A.1 Single-file motion
Eight sets of data from the complete experiment were used for empirical validation. Detailed information is presented in Tab.2. To avoid interference from external factors, 1000 frames of edge data were excluded. The middle segment of the data was sampled at a frequency of 2.5 Hz for analysis.
Table 2: Runs with number of candidates and global density.
Run name Number of candidates Global density (ped/m)
Young-10 10 0.39
Young-15 15 0.58
Young-20 20 0.78
Young-25 25 0.97
Young-30 30 1.17
Young-35 35 1.36
Young-40 40 1.56
Young-45 45 1.75
A.2 Unidirectional flow
The specific details of the unidirectional pedestrian flow experiment are presented in Tab.3. Similar to the simulation analysis, data from frames 501-600 (corresponding to 4 s) were collected at a sampling frequency of 25 Hz for the purpose of empirical validation.
@@ A.2 Unidirectional flow
Run name Walking direction Width b1 (m) Width b2 (m) Number of candidates

### v4

@@ 5.3 Discussion
Figure 19: Interaction graph of crowd: the topology and corresponding efficiency of the network govern the collective properties of the crowd. In subfigure (b), $\psi$ denotes the phase difference function. For further explanation of the space-speed time delay ($\delta$), please refer to Wang et al. (2024).
@@
In sections 5.1 and 5.2, we discussed the "catfish effect" and the process of crowd collapse, and provided an explanation based on the processes of critical transition. From the simulation results, we identified several densities of tipping point for crowd transitions, specifically $\rho=$ 9.8 and 10.4 ped/m2. These values are applicable only within the context of this simulation. Human body size varies significantly by gender, age, and other physiological factors, density value in the context of pedestrian crowd is an insufficient metric. More importantly, it is essential to understand the critical conditions under which individuals lose control of their motion due to contact. All insights in section 5 were derived retrospectively from the perspective of simulation with inherently limitations and potential errors. Based on the presented results, several issues need to be discussed in this section.
• Network dynamics perspective on crowd critical transition.
The network mapping of crowd interactions facilitates the analysis of critical transition processes from the perspective of network dynamics, thereby enabling a more concise understanding. As illustrated in Fig. 19(a), in non-contact scenarios, the anticipation and reaction behaviors of pedestrians can be characterized using mappings of arc connections (feedback mechanism) and edge connections (feedforward mechanism). The interactions among dense crowds correspond to structured networks. Such relationships can be extended to broader contexts and incorporated into a unified framework encompassing diverse collective systems, as demonstrated in Fig. 19(b). The topological structure of interaction graphs governs the dynamical evolution of crowds, where the connectivity influences the system’s response to conditional variations. In networks with incomplete connectivity that result in modular structures (characterized by low crowd density where pedestrians maintain certain distances), the system often exhibits a gradual adaptive capacity to changes. In contrast, within structured networks (featuring high crowd density with direct physical contact among pedestrians), individual disturbances are amplified through synergistic effects among linked nodes, ultimately driving the crowd system toward abrupt collapse under critical perturbations (critical transition) (Scheffer et al., 2012).
As illustrated in Fig. 19(c), in our simulation, the network undergoes a gradual transition from a sparse configuration to a structured architecture as density increases. During this process, the $\delta$ between nodes progressively converges to zero, leading to enhanced response efficiency in the crowd system. Within crowd contexts, this represents a dangerous property. When the $\delta\to 0$, it indicates the onset of mutual compression within the crowd, manifested through increased velocity coherence and correlation length. Should this trend continue to develop, the occurrence of a critical transition at the tipping point becomes inevitable.
• Whether dynamic (or aggressive) pedestrians would affect the process of crowd collapse.
@@
Potential intentional pushing was reported during the Itaewon crowd crush incident (Kim & Kim, 2022). Similar aggressive behaviors, as manifestations of crowd heterogeneity, require attention in simulations. In our simulation, introduction of dynamic pedestrians significantly destabilizes the state of the crowd. From the perspectives of mobility (average normalized speed) and symmetry (order parameter), each dynamic pedestrian acts like a microscopic magnet that induces local polarization in the crowd, as demonstrated in section 5.1. This phenomenon is characterized by an increase in observed dynamism and a concomitant reduction in symmetry, resulting in pronounced fluctuations in both velocity and density within localized regions. From the perspective of network dynamics, dynamic pedestrians can be analogized to active nodes. In a highly responsive structured network, such active perturbation generators will inevitably lead to a degradation of robustness.
@@ 6 Advantages and limitations
We have conducted extensive empirical validation and testing of the model. In this part, The most fundamental properties of the model will be discussed: how efficient the model is and under what circumstances the model may become ineffective.
@@ 6.1 Property of linear time complexity
 for $\textit{round}=1$ to 22 do
    if round == 2 and $\textit{step\_length}\bmod 2==1$ and $D_{ij}<(\textit{step\_length}-1)\cdot\ell/2$ then
@@
        Update location_grid $\mathrel{+}=\textit{direction\_search}$
@@
       case (0,1)(0,-1) do
          direction_search = (1,0)(-1,0)
@@
Benchmark tests were conducted on a laptop (Intel Core i7-11800H CPU, 16 GB memory) to evaluate simulation performance. The simulation tested five configurations for structured basic units: an adaptive length and four fixed lengths (0.5 m, 1 m, 1.5 m, and 2 m). In a 100 m $\times$ 100 m (10,000 m2) space, we randomly generated agents with equal increments (2,000 pedestrians per run) over 3,000 iterations. Fig.21 shows the results. The model clearly exhibited linear time complexity, with a statistically significant linear relationship between the basic unit length and the runtime. Comparisons across configurations revealed that larger unit sizes degraded computational efficiency, consistent with the linear fitting results. For adaptive unit length and fixed lengths of 0.5 m, 1 m, 1.5 m, and 2 m, the fitted slopes were 0.37, 0.33, 0.40, 0.53, and 0.69, respectively.
@@ 6.2 Limitations
We also explored the limitations of the model, which mainly include two aspects: Intractability in task related with decision process. As shown in Fig.22 (a) and (b), even simulating some local tactical-level behaviors, such as short-term route choice, is beyond the capability of the model, which is also a common limitation of most operational-level models. Distortion in response to static environments. In certain static environments, pedestrians may become trapped even when space is available. Within the framework of force-based models, there exist equilibrium positions in static environments that result in a balance of forces. These positions function analogously to black holes, attracting nearby pedestrians, as illustrated in Fig. 22(b).
@@ 7 Conclusions
Empirical validation of the model was conducted for scenarios involving single-file motion and unidirectional flow. In simulations, we examined the phase separation process in crowds, including lane formation, stripe formation, and cross-channel formation, typically regarded as the phenomena of pedestrian self-organization. By analyzing entropy decrease and steady-state characteristics, quantitative observations clearly reveal the self-organization process. Furthermore, through mass gathering simulations, we quantitatively demonstrated the "catfish effect", a theoretical concept that currently lacks empirical validation. In this context, we analyzed polarization and the process of symmetry variation in crowds.
Finally, focusing on crowd collapse, we explored potential mechanisms through time series and flow-field analysis. Applying critical transition theory, we characterized crowd collapse as a transition process quantified by the order parameter. As crowd density increases, the first transition (second-order) occurs when the crowd becomes tightly compressed. At this stage, pedestrian interactions shift from self-propelled constraints of fundamental diagrams to granular like dynamics. The second transition (first-order) occurs during crowd collapse. At this stage, further compression approaches the limit, elevating the order parameter. When the crowd reaches a critical state (characterized by high density and a persistently rising order parameter), collapse initiates.
Given the simple rules and evaluation based on fundamental parameters (e.g., average normalized speed, order parameter), assessments of crowd dynamics can be intuitively derived, thus providing valuable insights and inspiration. However, since many findings are derived from hindsight of simulations, their inherent limitations should not be overlooked. Furthermore, this model builds a bridge between traffic and pedestrian modeling on the basis of force-based methods.
@@
Code and data can be found at: https://drive.google.com/drive/folders/1gdkYmaUhcPgWp4i0Cffh2bph1fwvdQNp (Google Drive) and a video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
@@ Appendix A Experiments
The empirical validation in section 3 was conducted using data from experiment of unidirectional flow. The datasets are available at: https://ped.fz-juelich.de/da/doku.php?id=start#data_section (Pedestrian Dynamics Data Archive). Detailed experimental configurations and procedural specifics can be found in the works of Cao et al. (2017), as referenced. The specific details of the unidirectional pedestrian flow experiment are presented in Tab.2. Similar to the simulation analysis, data from frames 501-600 (corresponding to 4 s) were collected at a sampling frequency of 25 Hz for the purpose of empirical validation.
@@
Run name Direction Width b1 (m) Width b2 (m) Number of candidates